OpenStudy (anonymous):

write the slope intercept form of the equation of the line described.through:(3,3) parallel to y=1/4x+4

OpenStudy (anonymous):

parallel lines have same slope so use m=1/4 and (3,3) and solve for b just like before

OpenStudy (anonymous):

parallel means same slope so you would use 1/4 plus (3,3) in to y = mx + b using 1/4 to get your b

OpenStudy (anonymous):

y=1/4x+9/4

OpenStudy (anonymous):

exactly what i just got!

OpenStudy (anonymous):

AWESOME!